11 11 TEX L^{A}TEX

Linux Meta-FAQ

Michael K. Johnson <johnsonm@redhat.com> v4.7, 25 October 1997

This is the Meta-FAQ for Linux. It is mainly a list of valuable sources of information. Check these sources out if you want to learn more about Linux, or have problems and need help.

1 Introduction

What is Linux?

Linux is an independent implementation of the POSIX operating system specification, with SYSV and BSD extensions, that has been written entirely from scratch (this means it looks and acts just like Unix). It has no proprietary code in it. Linux is freely distributable under the GNU General Public License. Linux works on IBM PC compatibles with an ISA or EISA bus (including local bus variants VLB and PCI) and a 386 or higher processor. Some Amiga and Atari computers with MMU’s are also supported. This means 68020 with an external MMU, 68030, 68040, or 68060. Support for the Digital Alpha is now stable. Red Hat and Craftworks have Alpha distributions of Linux. Support for Sparc is stable, and Red Hat Linux is available for Sparc. Support for PowerPC is in development for multiple platforms, including Nubus and PCI Macintosh, Motorola Powerstack, IBM 830 and 850, and other platforms. Support for ARM, StrongARM, and MIPS is in various stages of completion, but don’t hold your breath. Read comp.os.linux.announce instead. See the Linux INFO-SHEET for more technical information on these ports, and the Hardware Compatibility HOWTO for more exact hardware requirements. The Linux kernel is written by Linus Torvalds <torvalds@transmeta.com> and other volunteers. Most of the programs running under Linux are generic Unix freeware, many of them from the GNU project.

The Linux INFO-SHEET More specific technical information on Linux. Includes pointers to information on the various ports, a feature list, information about how to get Linux, and more.

The Linux HOWTO’s These are somewhat like FAQ’s, but instead of answering common questions, they explain how to do common tasks, like ordering a release of Linux, setting up print services under Linux, setting up a

basic UUCP feed, etc. See <http://sunsite.unc.edu/LDP/HOWTO/HOWTO-INDEX.html> or <ftp:

//sunsite.unc.edu/pub/Linux/docs/HOWTO/> for the definitive versions of all the HOWTO’s. Other sites with up-to-date copies of the HOWTOs are ftp.cc.gatech.edu and tsx-11.mit.edu. In addition, there are many short, free-form documents called "mini-HOWTOs". These documents cover very specific subjects, such as BogoMIPS or Color-ls. These are available at <ftp://sunsite.unc.

edu/pub/Linux/docs/HOWTO/mini/> and at <http://sunsite.unc.edu/LDP/HOWTO/HOWTO-INDEX.
html>.

Linux newsgroups There are several Usenet newsgroups for Linux. It is a good idea to follow at least comp.os.linux.announce if you use Linux. comp.os.linux.announce is moderated by Lars Wirzenius. To make submissions to the newsgroup, send mail to linux-announce@news.ornl.gov. You may direct questions about comp.os.linux.announce to Lars Wirzenius <wirzeniu@iki.fi> The newsgroup comp.os.linux.announce is a moderated newsgroup for announcements about Linux (new programs, bug fixes, etc). The newsgroup comp.os.linux.answers is a moderated newsgroup to which the Linux FAQ, HOWTO documents, and other documentation postings are made. The newsgroup comp.os.linux.setup is an unmoderated newsgroup for discussion of issues and problems involved in setting up Linux systems. The newsgroup comp.os.linux.admin is an unmoderated newsgroup for discussion of administration of Linux systems. The newsgroup comp.os.linux.development.system is an unmoderated newsgroup specifically for discussion of Linux kernel development. The only application development questions that should be discussed here are those that are intimately associated with the kernel. All other development questions are probably generic Unix development questions and should be directed to a comp.unix group instead, unless they are very Linux-specific applications questions, in which case they should be directed at comp.os.linux.development.apps. The newsgroup comp.os.linux.development.apps is an unmoderated newsgroup specifically for discussion of Linux-related applications development. It is not for discussion of where to get applications for Linux, nor a discussion forum for those who would like to see applications for Linux. The newsgroup comp.os.linux.hardware is for Linux-specific hardware questions. The newsgroup comp.os.linux.networking is for Linux-specific networking development and setup questions.

The newsgroup comp.os.linux.x is for Linux-specific X Windows questions. The newsgroup comp.os.linux.misc is an unmoderated newsgroup for any Linux discussion that doesn’t belong anywhere else. In general, do not crosspost between the Linux newsgroups. The only crossposting that is appropriate is an occasional posting between one unmoderated group and comp.os.linux.announce. The whole point of splitting the old comp.os.linux group into many groups was to reduce traffic in each. Those that do not follow this rule will be flamed without mercy...

Other newsgroups Do not assume that all your questions are appropriate for a Linux newsgroup just because you are running Linux. Is your question really about shell programming under any unix or unix clone? Then ask in comp.unix.shell. Is it about GNU Emacs? Then try asking in gnu.emacs.help. Also, if you don’t know another group to ask in, but think there might be, politely ask in your post if there is another group that would be more appropriate for your question. At least the groups comp.unix.{questions,shell,programming,bsd,admin} and comp.windows.x.i386unix should be useful for a Linux user.

The World-Wide Web Greg Hankins <gregh@cc.gatech.edu> maintains the home WWW page for the Linux project. The

URL is <http://sunsite.unc.edu/LDP/>

Linux Journal A magazine called Linux Journal was launched several years ago. It includes articles intended for almost all skill levels, and is intended to be helpful to all Linux users. Subscriptions are $22 in the U.S., $27 in Canada and Mexico, and $32 elsewhere around the world, all payable in U.S. funds. Subscription inquiries can be sent via email to subs@ssc.com or faxed to (U.S.) 1-206-782-7191 or mailed to Linux Journal, PO Box 85867, Seattle, WA 98145-1867 USA. SSC has a PGP public key if you wish to send your credit card number via encrypted email: finger info@ssc.com

The Linux Software Map Information on free software available for Linux can be found in the Linux Software Map, which can

be found at <http://www.execpc.com/~lsm/>

2 Getting Linux

2.1 Linux FTP sites

A more complete list of Linux FTP sites is in the Linux INFO-SHEET, which can always be found at

<http://sunsite.unc.edu/LDP/HOWTO/INFO-SHEET.html> The most important sites are listed here; please

see the INFO-SHEET for a site nearer to you (there are many mirrors).

textual name             numeric addr    Linux directory
=======================  ==============  ===============
tsx-11.mit.edu           18.86.0.44      /pub/linux
sunsite.unc.edu          152.2.22.81     /pub/Linux
ftp.kernel.org           206.184.214.34  /pub/linux

These sites are the main “home” sites for Linux where most uploads take place. There are many mirror sites; please use the closest (network-wise) site to you.

2.2 Linux on physical media

Linux is distributed on physical media, mainly CD-ROM, by several commercial vendors. Please read the distribution HOWTO, posted regularily to comp.os.linux.announce, and available at <http://sunsite.

unc.edu/LDP/HOWTO/Distribution-HOWTO.html>

2.3 AFS

Linux is available over AFS by mounting the volume project.linux from sipb.mit.edu

2.4 Commercial networks

Compu$erve has some Linux archives.

2.5 Mailservers and such

Sunsite offers ftp-mail service — mail <ftpmail@sunsite.unc.edu>.

3 Linux distributions

Linux is distributed by its author only as a kernel. Other people have put together “distributions” that pair the Linux kernel with utilities and application software to make a complete working package.

There are several distributions of Linux, which are available at various sites. Sunsite mirrors many

of the distributions at <ftp://sunsite.unc.edu/pub/Linux/distributions/>. The most commonly-

recommended freely-available distributions are Red Hat <http://www.redhat.com> and Debian <http: //www.debian.org>. These are available for free over the internet, and are also sold on CD-ROM.

There are other distributions of Linux as well. Most commercial distributors of Linux advertise in Linux Journal.

4 Linux mailing-lists

Used mostly for discussion between developers of new features and testers of pre-release versions. See addresses in the FAQ. Send mail to majordomo@vger.rutgers.edu with the single word help in the body of the message , and you will get mail explaining how to subscribe to the many Linux mailing lists there. Save this mail, as it tells you how to unsubscribe from the lists, and if you post annoying messages to the list complaining about not being able to get off the list (because you didn’t follow instructions and save the mail telling you how to unsubscribe), you will likely be flamed for wasting international bandwidth and money.

5 Documentation for various programs

Many programs come with some sort of documentation, often in a file called README or something similar. It is a VERY good idea to read them with care. It is boring to see (and answer) questions that are answered in the documentation. Most programs also have “man pages”; use the command man programname to get documentation on a program named programname. To get help using the man program, use man man.

Most distributions put other documentation about programs in the directory /usr/doc/; your distribution should include documentation on how to access that documentation.

6 More Documentation

The Linux Documentation Project is working on a lot of documentation. Already, over 3000 pages of bookstyle documentation has been released to the general public, and another 2000 or so printed pages of man pages have also been released, with more to follow. Check <http://sunsite.unc.edu/LDP/> for documents written by the LDP.

7 Keeping track of current releases

Important new releases, programs, and ports are usually announced in comp.os.linux.announce.

8 This Document

The latest version of this document should always be available from <http://sunsite.unc.edu/LDP/

HOWTO/META-FAQ.html>

9 Legalese

Trademarks are owned by their owners. Satisfaction not guaranteed. No warranties about this document. Void where prohibited.

The content of this document is placed in the public domain, but if you quote it, please be polite and attribute your source.

Lars Wirzenius <wirzeniu@iki.fi> wrote the first version of this document; it is now maintained by Michael K. Johnson <johnsonm@redhat.com>. Mail me if you have any questions about this document.